OpenStudy (anonymous):

Solve the following system using substitution or elimination: The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x.

hero (hero):

2x + y = 46 y = x + 4

OpenStudy (anonymous):

y=x+4, 2x+y = 46 are the two simultaneous equations that you can derive from the information. Next use your favourite method of solving them (out of elimination or substitution).

OpenStudy (anonymous):

x=14 y=18

OpenStudy (anonymous):

2x + (x+4) = 46 ---> 3x = 42 ---> x=14

OpenStudy (anonymous):

substitution

hero (hero):

Okay, I'll use my favorite method: 2x + y = 46 y = x + 4 y = 46 - 2x y = y x + 4 = 46 - 2x 3x = 42 x = 42/3 x = 14 y = 18

OpenStudy (anonymous):

then put x=14 into either of your starting equations to get y=18